Suppose the two glass plates in are two microscope
Answer #1

The thicknessof the fringe \beta =\frac{L\lambda /2}{t}

where L is length of the plate

t is thickness of the paper

the fringe width \beta =\frac{0.12x250x10^{-9}}{0.022x10^{-3}}

\Rightarrow \beta =\frac{30}{0.022}x10^{-6}= 1.364x10^{-3} m

if thickness is increased, the denominator becomes twice, hence fringe width becomes half,

  \Rightarrow \beta' =0.681x10^{-3}m